Consider a bond which pays 8% semiannually and has 8 years to maturity. The market requires an interest rate of 8% on bonds of this risk. What is this bond's price? (Assume the Face Value of the bond is $1000)

a.$530.58

b.$891.62

c.$893.30

d.$3129.17
Answer #1

Information provided:

Face value= future value= $1,000

Time= 8 years*2= 16 semi-annual periods

Coupon rate = 8%/2= 4%

Coupon payment= 0.04*1,000= $40

Interest rate= 8%/2= 4%

The price of the bond is calculated by computing the present value.

Enter the below in a financial calculator the present value:

FV= 1,000

N= 16

PMT= 40

I/Y= 4

Press the CPT key and PV the present value.

The value obtained is 1,000.

Therefore, the price of the bond is $1,000.

In case of any query, kindly comment on the solution